You have a closed pipe which resonates it’s first harmonic when the pipe is .23m long.  What is the wavelength and frequency of the tuning form at 343m/s?


1
Expert's answer
2021-01-14T12:51:29-0500

"\frac{\lambda}{2}=0.23\to \lambda=0.46(m)" . Answer


"\nu=\frac{v}{\lambda}=\frac{343}{0.46}=745.6(Hz)" . Answer